# Line-of-sight shear in SLACS strong lenses

### v1

@@ 4.3 Adding boxyness and diskyness
where $\theta$ and $\varphi$ are the radial and angular coordinates of $\bm{\theta}$, $m$ is the order of the multipole, $a_{m}$ is its strength and $\varphi_{m}$ is the orientation angle of its main axis (Keeton et al., 2003; Xu et al., 2015). In the case of the octupole, $m=4$, and the shape is called ‘boxy’ when $\varphi_{4}=\frac{\pi}{4}$ and ‘disky’ when $\varphi_{4}=0$. To investigate whether boxyness or diskyness can affect LOS shear measurements, we further modelled our 50 SLACS lenses with the EPL_BOXYDISKY profile in lenstronomy, which adds a pure octupolar distortion on top of the elliptical power law profile (Van de Vyvere et al., 2022). We kept the other components of the lens and light models the same.
We find that, in general, the addition of the octupole to the lens mass potential does not lead to significant reductions in the LOS shear measurement; in fact, the shear measurements are typically the same or of greater magnitude than the measurements made using the EPL and minimal LOS shear model alone. The octupole strength, $a_{4}$, tends to favour large values, leading to a lens convergence with a highly octupolar shape. We show an example of the reconstruction of lens SDSSJ1020+1122 in Figure 9, using our standard EPL and minimal LOS shear model (top row) and with the addition of the octupole component (bottom row). This figure shows the reconstruction of the image, the residual difference between the reconstruction and the data, the reconstructed source, the convergence map and finally the joint posterior distribution of the LOS shear components.
Figure 9: The fit of lens SDSSJ1020+1122 with the EPL + minimal LOS shear model (top row) and the EPL + minimal LOS shear + octupole model (bottom row). From left to right, the panels show the reconstruction of the image, the residual difference between the image data and the reconstruction, the source reconstruction, the convergence map and the joint posterior distribution of the components of the LOS shear, $\gamma_{1}$ and $\gamma_{2}$. The dark and light shaded areas are respectively the $1\sigma$ and $2\sigma$ confidence intervals.
The effect of the octupole is not particularly evident in the reconstruction of the image, nor in the residuals or the value of the reduced $\chi^{2}$. However, it clearly leads to a strongly octupolar convergence and a very different source reconstruction to the case of the minimal model. The increased uncertainty on the LOS shear components can also be seen in the larger $1\sigma$ and $2\sigma$ regions of the posterior.
In Figure 10, we show the distribution of the measured LOS shear in both the minimal model and the minimal model plus octupole, for the same lenses in each case. This plot provides evidence that the inclusion of an octupole in the elliptical power law does not lead to LOS shear measurements which are more in line with expectations from weak lensing, i.e. at or around the percent level, for this subset of strong lenses. From this we conclude that, if shear measurements that are in agreement with weak lensing are the goal, boxyness or diskyness should only be included in the lens model if physically well-motivated for the specific lens in question. This conclusion is supported by the comparison shown in Figure 8 of the distributions of the individual shear components from simulations and from the observed lenses, potentially pointing to a selection effect wherein the lenses studied in this and other works are those that, by chance, have larger shears than expected from weak lensing, rather than there being an underlying physical cause for the difference. We leave further consideration of this for a future work.
Figure 10: Distribution of the LOS shear measurements in the EPL + minimal LOS shear model (solid blue) and the EPL + minimal LOS shear + octupole model (unfilled orange).
4.4 Does lens light trace lens mass?
The dolphin pipeline allows the user to apply a prior which joins the ellipticity of the lens mass with the lens light, restricting the differences between the axis ratios and position angles of the two components. This prior appears to be physically well-motivated, as it is expected that the lens light traces the underlying lens mass, within some reasonable uncertainty. However, if there are strong foreground effects, the lens light will experience significant foreground lensing, meaning that it may well no longer trace the lens mass within the prior constraints. To explore this scenario, we applied the aforementioned prior in our modelling pipeline, fitting the lenses with the minimal shear model, and requiring that the lens light ellipticity of the first Sérsic profile and lens mass ellipticity position angles be within 15 degrees of each other, and that the axis ratio of the lens light and lens mass ellipticity should be the same.
In Figure 11, we show the one and two dimensional marginalised posterior distributions on the lens mass ellipticity $e^{\rm mass}$, the ellipticity of the first Sérsic lens light profile $e^{\rm light}$, along with the minimal model shear parameters, $\gamma_{\rm od}$ and $\gamma_{\rm LOS}$ that result from fitting the lens SDSSJ0330-0020 with our lens model, with (red contours) and without (blue contours) the ellipticity prior described above. Furthermore, we also show the posterior distributions as obtained when fitting this lens with the no foreground shear model (purple contours).
@@ 4.4 Does lens light trace lens mass?
From this figure, we can see that the prior acts to confine the joint posterior of the lens mass and light ellipticity components to a small region of parameter space, as intended. However, due to the well-known degeneracy between lens ellipticity and foreground shear, this has the knock-on effect of also restricting the posteriors of the foreground shear parameters. The end result is that the LOS shear constraint for this lens changes when this prior is included, from $\gamma_{\rm LOS}=0.03\pm 0.022$ without the prior to $\gamma_{\rm LOS}=0.05\pm 0.071$ with the prior, due to the too-stringent restriction of the lens ellipticity.
Whilst these two LOS shear measurements are still consistent at $1\sigma$, this may not always be true, particularly in the case of large foreground shears where we expect the effect to be greater. This conclusion is borne out by the fact that the constraints on these parameters in the no foreground shear model are virtually identical to the constraints obtained in the minimal model when this prior is applied. We therefore argue that such priors should always be avoided when shear measurements are attempted. We have shown a single example here but this effect can be seen in all the lenses where the minimal model was successful.
@@ 5 Conclusions
In this work, we modelled a selection of 50 of the SLACS strong lenses, aiming to robustly measure the LOS shear for the first time. We used a lens model which consisted of an elliptical power law profile for the main deflector, two Sérsic profiles for the lens light, and a Sérsic plus shapelets for the source. On top of this, we modelled the LOS shear using the minimal model of Fleury et al. (2021). We found that, of these 50 lenses, only 18 were successfully fit with this model, with the majority of rejected fits being due to strong residuals between the best-fit model and the data, indicative of an overall poor lens model. We found that the distribution of our LOS shear magnitude measurements is consistent with those in the recent literature, obtained using the usual ‘external’ or ‘residual’ shear model.
We examined the effect of neglecting the post-Born corrections to the main deflector due to foreground shear and the LOS rotation in our model, re-fitting the same lenses with these parameters fixed to zero. We found that the degeneracy present between the lens mass ellipticity and the foreground shear is propagated to the LOS shear, and for certain individual lenses this leads to significantly different LOS shear measurements. However, considering the sample as a whole, this bias appears to average out for small shear magnitudes.
We also fitted our lenses with a further degree of freedom in the lens mass model, allowing for octupolar distortions, i.e. boxyness and diskyness. We found that this leads to larger LOS shear magnitudes in general, and in some specific cases, significantly different constraints on the LOS shear parameters to those found using the minimal model.
Lastly, we investigated the use of a prior which joins the lens mass and lens light ellipticities. We found that this prior has the same effect on LOS shear measurements as the removal of foreground shear from the model, with shear measurements becoming biased compared to the minimal model. This is due to the prior introducing extremely tight constraints on the lens mass ellipticity, which propagates to the LOS shear.
@@
It is also clear that the question of whether the total magnitude of the LOS shear parameter is produced by line-of-sight effects remains open. We found LOS shear measurements in good agreement with external shear measurements from the recent literature, which are known to be larger than weak lensing shear estimated around strong lenses (Etherington et al., 2024), as well as those from simulations, as discussed above. This is the key point we wish to emphasise: given the agreement between our findings and those of previous works, we conclude that the lines of sight of the SLACS strong lenses studied in this work are well-characterised and modelled. It is the modelling of the main deflector to which attention must now be devoted, if the cosmological shear signal is to be disentangled from the contribution of the lens. However, our results also show that adding boxyness and diskyness to the lens mass model is not the panacea that might have been expected, at least for the strong lenses modelled in this work. Whether any further complexity can be added to mass models that is still physically well-motivated remains to be seen.
@@
In conclusion, the task of uniformly modelling a very large catalogue of strong lenses and obtaining accurate shear measurements for the purposes of cosmology will clearly not be achieved with current data or modelling techniques. Lens model constraints may be improved using image data in which the lensed source light is more easily separable from the lens light; this may be seen in James Webb Space Telescope (JWST) imaging of high-redshift lensed sources e.g. Mercier et al. (2024) – since JWST observes in the near and mid-infrared, it has a much higher signal-to-noise ratio for redder sources compared to e.g. HST. Furthermore, more work must be done to build lens models which accurately match the physical complexity of strong lens galaxies.
Acknowledgements.
@@
We are very grateful to Théo Duboscq, Pierre Fleury and Giacomo Queirolo for useful discussions throughout the preparation of this work and for their insightful comments on the manuscript.

### v2

@@ 4.3 Shear does not compensate for unmodelled octupoles
where $\theta$ and $\varphi$ are the radial and angular coordinates of $\bm{\theta}$, $m$ is the order of the multipole, $a_{m}$ is its strength and $\varphi_{m}$ is the orientation angle of its main axis (Keeton:2002qt; Xu:2014dda). In the case of the octupole, $m=4$, and the shape is called ‘boxy’ when $\varphi_{4}=\frac{\pi}{4}$ and ‘disky’ when $\varphi_{4}=0$. To investigate the effect the inclusion of an octupole has on the LOS shear, we re-model our sample of lenses using the EPL_BOXYDISKY profile in lenstronomy, which adds a pure octupolar distortion on top of the elliptical power law profile (VandeVyvere2021). The source models for each lens are kept the same. We use the minimal LOS shear model.
Figure 7: This figure shows the measured value of $|\gamma_{\rm LOS}|$ with the $m_{4}$ multipole included in the lens mass model. The black dotted line shows the overall mean expected value of $|\gamma_{\rm LOS}|$, computed from the RayGalGroup simulations, as described in the main text, with the associated 1, 3 and $5\sigma$ uncertainties around this mean given by the grey shaded bands. The measured values are coloured by the change in the tension between the minimal model and the minimal + octupole model.
Figure 7 shows the measurement of $|\gamma_{\rm LOS}|$ in each lens with the addition of the octupole to the mass model. The systems are ordered by increasing deflector redshift on the $x$-axis. The black dotted line and grey bands represent the mean and $1$, 33 and $5\sigma$ uncertainties on $|\gamma_{\rm LOS}|$ inferred from the RayGalGroup simulations, as described above. In contrast to Figure 6, we plot the overall mean and uncertainties rather than those quantities computed for each lens, hence the absence of the redshift-dependent slope seen in that previous figure.
Furthermore, the measured points in this plot are coloured according to the change in the tension statistic compared to the result using the minimal LOS model without the additional octupole in the lens mass. All values are listed in Table 1. Points coloured in orange are in greater tension with the mean of the simulations compared to the measurement in the minimal model, whilst points coloured in blue are in lesser tension. We note that for SDSSJ0959+0410 and SDSSJ2303+1422, either one or both of the LOS shear components $\gamma_{1}$ and $\gamma_{2}$ were unconverged in the octupole model; we do not include these lenses on this plot.
@@
Figure 8: This figure shows the residual difference between model and data when modelling SDSSJ1204+0358, SDSS0737+3216 and SDSSJ1531-0105 using the minimal model (left panel) and with the inclusion of an octupole (middle panel). The right-hand panel shows the 1D marginalised posterior distribution of the LOS shear magnitude $|\gamma_{\rm LOS}|$ in the minimal model (blue) and with the inclusion of an octupole (yellow).
The final remaining three lenses, SDSSJ1204+0385, SDSSJ0737+3216 and SDSSJ1531-0105, show changes in tension, $\Delta\sigma$, of 2.2-2.2, 3.4-3.4 and 4.5-4.5 respectively. In Figure 8, we show the residual difference between data and model for the reconstructions of these lenses in the minimal model and with the octupole added to the mass model. We also show the 1D marginalised posterior distributions for the LOS shear magnitude, $|\gamma_{\rm LOS}|$, inferred in these two models (minimal model in blue, model with additional octupole in yellow).
For lens SDSSJ1204+0385, we can see two things. Firstly, with the octupole included in the model, the residuals between the data and model are significantly increased. Secondly, we can see that the apparent reduction in tension between the $|\gamma_{\rm LOS}|$ value inferred in the octupole model and the expectation from simulations for this lens is driven by a large increase in the uncertainty on the $|\gamma_{\rm LOS}|$ measurement; the error bars are over ten times larger in the octupole model than in the minimal model.
Another thing to note is that our baseline model is nested within the octupole model, meaning that, if the baseline model was the true solution to the lens equation for this system, we should expect that the octupole model would recover the baseline model via the octupole strength $|a_{4}|$ going to zero. For this lens, we find that $|a_{4}|=-0.02^{+0.07}_{-0.04}$, which is consistent with zero at $1\sigma$. Nevertheless, a different solution to the lens equation is clearly found, with the combination of a more concentrated source and shallower EPL slope than in the baseline model being recovered. This may point to the instability of the solutions found by either or both models.
For lens SDSSJ0737+3216, the two models give virtually identical fits to the data. Again, the reduction in tension with the simulated data is brought about by the slight loss of precision on $|\gamma_{\rm LOS}|$ when the octupole is included in the lens mass model. However, with equivalent fits and identical reduced $\chi^{2}$ values of $0.83$, Ockham’s razor implies that, for this particular lens, we should prefer the model with the fewest parameters [^3] The most principled way to determine the best-fitting model for each lens would be to carry out a model comparison using the Bayesian evidence, which may be obtained by using nested sampling to compute the normalising factor in Bayes’ theorem, e.g. (Skilling2004; Handley2015a). An attempt at population-level lens model comparison would benefit from a larger sample of lenses than that used in this work. i.e. the minimal model without the octupole.
For lens SDSSJ1531-0105, the inclusion of the octupole serves to reduce the residuals between model and data present in the result using the minimal model, implying a better fit of the octupole model to the data. The $|\gamma_{\rm LOS}|$ value is shifted to a lower value, distinct from the result in the minimal model at just over $1\sigma$. This is the single case in our sample of lenses where the inclusion of the octupole both provides a better fit to the data than the minimal model and also reduces the LOS shear magnitude. From this we can conclude that unmodelled octupoles are not universally responsible for larger-than-expected shear magnitudes, but may be a necessary addition to a lens mass model in some specific cases.
Lastly, let us comment on the inferred octupole strength, $|a_{4}|$. In $35\%$ of the lenses studied, we recovered an octupole strength equal to or greater than $0.02$ i.e. $2\%$, whilst a further $45\%$ have octupole strengths greater than $5\%$. This is in contrast to the findings of various observational studies of early-type galaxies, which have consistently found octupolar distortions at the $1\%$ level (Hao2006; Pasquali2006; Goullaud2018).
The inclusion of the octupole was justified in this and previous works as an aspect of mass model complexity that is well-motivated by observations. Our finding that, in the majority of the lenses studied in this work, the octupole strength, along with the shear magnitude, is larger than expected, indicates that the $m_{4}$ multipole alone also does not capture the full complexity of strong lens galaxy mass profiles. We leave the exploration of other possibilities, such as the addition of an $m_{1}$ or $m_{3}$ multipole (Amvrosiadis2025; Lange2025), ellipticity gradients and isophotal twists (VandeVyvere:2022gqa), or LOS flexion (Duboscq:2024asf), to a future work.
4.4 Neglecting post-Born corrections affects LOS shear measurements
Hogg:2022ycw demonstrated that, in mock data, removing the post-Born correction to the potential of the main deflector which is induced by foreground shear (achieved by fixing $\gamma_{\rm od}=0$ in the parameter inference) had little effect on the inferred LOS shear magnitude. We now examine whether this finding holds true in real data; if it does, this would be useful from a parameter inference perspective, as it would reduce the dimensionality of the parameter space by two.
Figure 9: The one and two-dimensional joint marginalised posterior distributions for the lens mass ellipticity, foreground shear and LOS shear in lens SDSSJ1630+4520. The posteriors recovered in the minimal model are shown in blue and those recovered in the no foreground shear model in purple.
[…]
From this figure, we can see that when the foreground shear is excluded from the model, there are two consequences. Firstly, the inferred lens mass ellipticity is recovered with significantly higher precision than in the minimal model. Secondly, the posteriors of the LOS shear components are shifted with respect to those in the minimal model. In other words, with the enforced absence of the strong degeneracy with foreground shear, the lens mass ellipticity is recovered with artificially high precision. Globally, a slightly different best-fit lens model is therefore obtained, with a slightly different inferred LOS shear.
Figure 10: Left: $|\gamma_{\rm LOS}|$ measurements in the minimal model versus the no foreground shear model. Right: the same but restricted to show only $|\gamma_{\rm LOS}|<0.1$. The black solid line represents a linear ordinary least squares fit to the measurements, with the $1\sigma$ uncertainty shown by the grey shaded band. The $y=x$ line is shown by the grey dashed line.
We can explore whether this effect is systematic across our whole lens sample. In Figure 10, we show our measurements of $|\gamma_{\rm LOS}|$ in the minimal model on the $x$-axis plotted against the $|\gamma_{\rm LOS}|$ in the no foreground shear model on the $y$-axis. The left panel shows the full sample whilst the right panel restricts the view to values of the LOS shear magnitude less than 0.1, to improve the clarity of the cluster of measurements between 0 and 0.05. Furthermore, the black solid line and grey shaded bands show a linear fit [^4] To do this, we use the ordinary least squares fitting method provided by numpy.polyfit. However, least squares fitting becomes inaccurate when the independent variable has significant uncertainties, as is the case here (Fuller1987). We further choose not to weight the fit using the uncertainty on the dependent variable. The displayed uncertainty on each linear fit (shaded grey region) is therefore computed from the covariance matrix returned by numpy.polyfit based solely on the shear values, without their errors. and $1\sigma$ uncertainty, whilst the grey dashed line shows $y=x$.
This figure shows that for large shear values, $|\gamma_{\rm LOS}|>0.1$, the LOS shear magnitude measured in the no foreground shear model tends to be slightly smaller than what would be measured in the minimal model, whilst at small shear values (which themselves are likely to be more trustworthy), the inverse is true: the LOS shear magnitude tends to be slightly smaller in the minimal model than when the foreground shear is neglected.
From this we may conclude that the small speed enhancement engendered by the removal of the foreground shear parameters from the inference is not worth the possible bias this will induce on the inferred LOS shear magnitude; in other words, physical accuracy should not be traded for speed.
@@ 5 Conclusions
In this work, we modelled 23 of the SLACS strong lenses, aiming to robustly measure the LOS shear for the first time. We used a lens model which consisted of an elliptical power law profile for the main deflector, two Sérsic profiles for the lens light, and a Sérsic plus shapelets for the source. On top of this, we modelled the LOS shear using the minimal model of Fleury:2021tke. We found that the distribution of our LOS shear magnitude measurements is consistent with those in the recent literature, obtained using the usual ‘external’ or ‘residual’ shear model.
We examined the effect of neglecting the post-Born corrections to the main deflector due to foreground shear in our model, re-fitting the same lenses with these parameters fixed to zero. We found that this leads to small changes in the inferred LOS shear magnitude compared to the minimal model, with the differences becoming more significant when the inferred shear is large.
@@
We also fitted our lenses with a further degree of freedom in the lens mass model, allowing for octupolar distortions, i.e. boxyness and diskyness. We found that shear measurements are not generally reduced by the addition of this octupole, and only a single lens in our sample, SDSSJ1531-0105, was better fit by the octupole model than the minimal model. Nonetheless, measured shear values do in several cases show sensitivity to its inclusion.
@@
It is also clear that the question of whether the total magnitude of the LOS shear parameter is produced by line-of-sight effects remains open. We found LOS shear measurements in good agreement with external shear measurements from the recent literature, which are known to be larger than weak lensing shear estimated around strong lenses (Etherington:2023yyh), as well as those from simulations, as discussed above. This is the key point we wish to emphasise: given the agreement between our findings and those of previous works, we conclude that the shear contribution from the lines of sight of the SLACS strong lenses studied in this work are adequately captured by the lens model. It is the modelling of the main deflector to which attention must now be devoted, if the cosmological shear signal is to be disentangled from the contribution of the lens. However, our results also show that adding boxyness and diskyness to the lens mass model is not the panacea that might have been expected, at least for the strong lenses modelled in this work. Whether any further complexity can be added to mass models that is still physically well-motivated remains to be seen.
@@
In conclusion, the task of uniformly modelling a very large catalogue of strong lenses and obtaining accurate shear measurements for the purposes of cosmology will clearly not be achieved with current data or modelling techniques. Lens model constraints may be improved using image data in which the lensed source light is more easily separable from the lens light; this may be seen in James Webb Space Telescope (JWST) imaging of high-redshift lensed sources e.g. Mercier2024 – since JWST observes in the near and mid-infrared, it has a much higher signal-to-noise ratio for redder sources compared to e.g. HST. Furthermore, more work must be done to build lens models which accurately match the physical complexity of strong lens galaxies. The promise of shear measurements from strong lensing as a novel cosmological probe motivates the pursuit of advancements in models and modelling techniques, and progress continues on each of these fronts.
Acknowledgments
@@
We are very grateful to Théo Duboscq, Pierre Fleury and Giacomo Queirolo for useful discussions throughout the preparation of this work and for their insightful comments on the manuscript. We are also grateful to an anonymous referee whose comments led to a substantial improvement of our results.
NBH is supported by the research environment and infrastructure of the Handley Lab at the University of Cambridge. DJ acknowledges support by the First Rand Foundation, South Africa, and the Centre National de la Recherche Scientifique of France.
Data and software availability
The imaging data for the SLACS lenses studied in this work are publicly available. The lens modelling carried out in this work used the open-source dolphin package (Shajib:2025bho), with lenstronomy (Birrer:2018xgm; Birrer2021) as the modelling engine. The inputs and outputs of our modelling, including lens image cutouts, PSFs and MCMC chains, can be found at the following Zenodo repository: https://zenodo.org/records/17816315. A Jupyter notebook which reproduces the figures and tables in this work can be found at the following Github repository: https://github.com/nataliehogg/los_in_slacs.
The following colour-map packages were used in this work: cubehelix (Green2011), cmocean (Thyng2016) and cmasher (vanderVelden2020). NBH acknowledges the use of the AI coding agent Claude Code (Sonnet 4.5) for software development tasks. All AI-generated outputs were validated by NBH to ensure their accuracy.
Name Filter $n_{\rm max}$ $z_{\rm lens}$ $z_{\rm source}$ $|\gamma_{\rm LOS}|$ Tension $\sigma$] $\Delta\sigma$
SDSSJ0029-0055 F606W 6 0.227 0.931 $0.029^{+0.009}_{-0.013}$ 1.30   0.61
SDSSJ0037-0942 F606W 6 0.195 0.632 $0.078^{+0.015}_{-0.016}$ 3.88   0.26
SDSSJ0252+0039 F606W 6 0.280 0.982 $0.022^{+0.010}_{-0.010}$ 0.85 0.24-0.24
SDSSJ0330-0020 F606W 6 0.351 1.071 $0.049^{+0.013}_{-0.012}$ 1.99 1.14-1.14
SDSSJ0728+3835 F606W 16 0.206 0.688 $0.042^{+0.017}_{-0.013}$ 2.16   1.53
SDSSJ0737+3216 F [^5] W 6 0.322 0.581 $0.090^{+0.010}_{-0.010}$ 4.86 3.36-3.36
SDSSJ0903+4116 F606W 6 0.430 1.065 $0.052^{+0.008}_{-0.008}$ 1.98   3.46
SDSSJ0959+0410 F [^5] W 16 0.126 0.535 $0.040^{+0.009}_{-0.008}$ 3.34   n/a
SDSSJ1112+0826 F606W 6 0.273 0.629 $0.226^{+0.032}_{-0.030}$ 6.65 0.57-0.57
SDSSJ1204+0358 F606W 6 0.164 0.631 $0.051^{+0.004}_{-0.004}$ 4.89 2.15-2.15
SDSSJ1250+0523 F [^5] W 16 0.232 0.795 $0.025^{+0.005}_{-0.006}$ 1.45 0.15-0.15
SDSSJ1306+0600 F606W 6 0.173 0.472 $0.042^{+0.020}_{-0.021}$ 1.62 0.17-0.17
SDSSJ1313+4615 F606W 10 0.144 0.339 $0.024^{+0.010}_{-0.010}$ 1.62   3.32
SDSSJ1402+6321 F [^5] W 6 0.205 0.481 $0.008^{+0.010}_{-0.006}$ 0.14   0.25
SDSSJ1531-0105 F606W 6 0.160 0.744 $0.214^{+0.025}_{-0.025}$ 7.85 4.48-4.48
SDSSJ1621+3931 F606W 6 0.245 0.602 $0.028^{+0.009}_{-0.009}$ 1.39   5.48
SDSSJ1627-0053 F [^5] W 6 0.208 0.524 $0.023^{+0.004}_{-0.005}$ 1.41   0.84
SDSSJ1630+4520 F [^5] W 6 0.248 0.793 $0.046^{+0.005}_{-0.005}$ 3.29 0.34-0.34
SDSSJ1636+4707 F606W 16 0.228 0.674 $0.062^{+0.016}_{-0.032}$ 1.64 0.81-0.81
SDSSJ2238-0754 F [^5] W 16 0.137 0.713 $0.034^{+0.024}_{-0.028}$ 1.02 0.42-0.42
SDSSJ2300+0022 F [^5] W 16 0.229 0.464 $0.044^{+0.014}_{-0.014}$ 2.15 0.22-0.22
SDSSJ2303+1422 F [^5] W 6 0.155 0.464 $0.000^{+0.018}_{-0.000}$ 0.57   n/a
SDSSJ2343-0030 F606W 6 0.181 0.463 $0.081^{+0.013}_{-0.016}$ 4.17   1.73
Table 1: The SLACS strong lenses studied in this work. All redshifts are taken from Bolton2008. The $|\gamma_{\rm LOS}|$ values listed are those inferred in the minimal LOS shear model. The $\Delta\sigma$ values are the change in tension with the expectation from simulations when the octupole is added to the lens model.
Component Parameter Prior
$\gamma^{\rm EPL}$ $[1.3,2.8]$
$e_{1}$ [0.5,0.5][-0.5,0.5]
Lens mass $e_{2}$ [0.5,0.5][-0.5,0.5]
$x$ $0.5^{\prime\prime},0.5^{\prime\prime}]$
$y$ $0.5^{\prime\prime},0.5^{\prime\prime}]$
$R_{\rm{eff}}$ $[0.1,5.0]$
$e_{1}$ [0.5,0.5][-0.5,0.5]
Lens light $e_{2}$ [0.5,0.5][-0.5,0.5]
$x$ $0.5^{\prime\prime},0.5^{\prime\prime}]$
$y$ $0.5^{\prime\prime},0.5^{\prime\prime}]$
$\beta$ $\log_{10}[0.02^{\prime\prime},0.2^{\prime\prime}]$
$R_{\rm{eff}}$ $0.04^{\prime\prime},0.5^{\prime\prime}]$
$n_{\rm{S}}$ $[0.5,8.0]$
Source light $e_{1}$ [0.5,0.5][-0.5,0.5]
$e_{2}$ [0.5,0.5][-0.5,0.5]
$x$ $0.2^{\prime\prime},0.2^{\prime\prime}]$
$y$ $0.2^{\prime\prime},0.2^{\prime\prime}]$
$\gamma_{1}^{\rm od}$ [0.2,0.2][-0.2,0.2]
$\gamma_{2}^{\rm od}$ [0.2,0.2][-0.2,0.2]
Shear $\gamma_{1}^{\mathrm{LOS}}$ [0.5,0.5][-0.5,0.5]
$\gamma_{2}^{\mathrm{LOS}}$ [0.5,0.5][-0.5,0.5]
$\omega_{\mathrm{LOS}}$ [0.2,0.2][-0.2,0.2]
$a_{4}$ [0.1,0.1][-0.1,0.1]
Octupole $\phi_{4}$ $\pi,\pi]$
$x$ $0.5^{\prime\prime},0.5^{\prime\prime}]$
$y$ $0.5^{\prime\prime},0.5^{\prime\prime}]$